1. Suppose in year 2014, the amount of money in circulation is $400 billion, the price level is
120 and the real output is $100 billion.
a) What is the velocity of money?
b) Suppose in year 2015, real output became $102 billion, and the Fed increased the stock of
money to $420 billion. What is the inflation rate from 2014 to 2015? (
1
Expert's answer
2018-11-12T15:44:09-0500

a) MV = PY - Fisher`s formula

V = PY / M = 1,2 * 100 / 400 = 0,3 - velocity of money in 2014


b) in 2015 we see that all indicators have increased in 1,05, so we will find an inflition rate with GDP deflator:

GDP deflator = (nominal GDP / real GDP)* 100%

2014: GDP deflator = 120

2015: GDP deflator = 126

Inflation Rate = ((def2015 - def2014) / def2014) x 100 ;

Inflation rate = ((126 - 120 / 120) * 100 = 5%


So, inflation has increased in 5 %